OpenStudy (anonymous):

in the given function f(x)=3x+5y-30, what is the ordered pair for the y-intercept

OpenStudy (anonymous):

let x=0, calculate y y-int (0,y)

OpenStudy (anonymous):

to find y intercept set x=o and solve for y

OpenStudy (anonymous):

the 5y? or the f(x)

OpenStudy (anonymous):

f(0)

OpenStudy (anonymous):

f(x) that is the y?

OpenStudy (anonymous):

not sure of why your problem has f(x) and 5y...seems odd.. but nonetheless, y = 3x+5y-30 make x = 0 and solve for y

OpenStudy (anonymous):

no that is not correct

OpenStudy (anonymous):

y=3x+5y-30 ll be y intcpt x=0 let put 0 ll be y=3(0)+5y-30

OpenStudy (anonymous):

yes...dont forget about the y on the left side...per her problem

OpenStudy (anonymous):

then y-5y=-30 ll be -4y=-30

OpenStudy (anonymous):

-4y/-4=-30/-4

OpenStudy (anonymous):

correct...but please let the person who is asking the problem solve so that she understands.

OpenStudy (anonymous):

oky

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the answer was not in the choices

OpenStudy (anonymous):

i have been performing the formula that you have given

OpenStudy (anonymous):

a. (6,0) b. (0,6) c(10,0) d. (0,10)

OpenStudy (anonymous):

then confirm your problem....

OpenStudy (anonymous):

the answer is d

OpenStudy (anonymous):

but I don't know how get that answer by using a formula

OpenStudy (anonymous):

pls confirm your problem, based on what you asked that answer doesn't make sense to me

OpenStudy (anonymous):

I used trail and error method but we must show the solution on how we derived that answer

OpenStudy (anonymous):

hhmmm. JuanitaM that was the question already... in the given function f(x)=3x+5y-30, what is the ordered pair for the y-intercept

OpenStudy (anonymous):

and the choices are a. (6,0) b. (0,6) c(10,0) d. (0,10)

OpenStudy (kropot72):

How can it be correct to write f(x) = 3x + 5y -30 when the expression on the right hand side is a function of x and a function of y?